OpenStudy (hannahduhe):

given the point (3, -6) and the slope of 1/3, which of the following is the equation of the line? a.y=1/3x+7 b.y=1/3x+5 c.y=1/3x-7 d.y=-3x-7

TheSmartOne (thesmartone):

y = mx + b substitute the slope in place of m, and then plug in the point The point is in the form of (x, y) So, if the point is (1, 2), you plug in x = 1, and y = 2

TheSmartOne (thesmartone):

But, that's an example. The slope is 1/3 Replace that with m What is the x-coordinate in the point (3, -6) What is the y-coordinate?

OpenStudy (hannahduhe):

@TheSmartOne i know the first part of the equation is going to be y=1/3x but how do you find the next part

TheSmartOne (thesmartone):

y = (1/3)x + b Now, we need to plug in a value for x and y. What is the x-coordinate in the point (3, -6)? What is the y-coordinate of that point?

OpenStudy (hannahduhe):

is the equation b)y=1/3x+5?? @TheSmartOne

TheSmartOne (thesmartone):

No Answer the questions I asked you

OpenStudy (hannahduhe):

the x -coordinates of (3,-6) is 5 right @TheSmartOne

TheSmartOne (thesmartone):

\(\Large\bf (3, -6)\) \(\Large\bf (x,~ y)\) What is the number in place of x? What is the number in place of y?

OpenStudy (hannahduhe):

so x=3 and y=-6 right @TheSmartOne

TheSmartOne (thesmartone):

yes, plug that into the equation and solve for b

OpenStudy (hannahduhe):

@TheSmartOne how do you solve for b??

TheSmartOne (thesmartone):

y = (1/3)x + b -6 = (1/3)(3) + b What is (1/3) * 3 = ?

OpenStudy (hannahduhe):

1/3*3=1 @TheSmartOne

TheSmartOne (thesmartone):

- 6 = 1 + b Subtract 1 on both sides b = -6 - 1 b = ??

OpenStudy (hannahduhe):

b=-7 @TheSmartOne

OpenStudy (hannahduhe):

@TheSmartOne so the answer is c.

TheSmartOne (thesmartone):

@hannahduhe Yes

OpenStudy (hannahduhe):

@TheSmartOne thanks for all your help i really appreciate it

TheSmartOne (thesmartone):

Anytime! :)